OpenStudy (anonymous):

Let f(x) = x**6. f(1) =?

OpenStudy (anonymous):

I mean x^6

OpenStudy (xishem):

\[f(x)=x^6\]\[f(1)=(1)^6=1\]

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

wait if you got f(1)= 1 then can you show me how At x = 1, the slope of the tangent line to the graph =

OpenStudy (anonymous):

basically find the slope I think

OpenStudy (xishem):

Yep. If you take the derivative of the general function, you will get a function that outputs the slope of the tangent line at input, x... Given that...\[f(x)=x^6\]Then...\[f'(x)=6x^5\]The slope at x = 1 is found by plugging 1 into the derivative of the function f(x)...\[f'(1)=6(1)^5=6\]